# Copper delocalization leads to ultralow thermal conductivity in chalcohalide CuBiSeCl2

### v1

@@ Introduction
Mixed-anion compounds, such as halide-chalcogenides, have recently gained attention as promising candidates for thermoelectric and photovoltaic applications due to their ability to achieve low thermal conductivity while maintaining favorable electronic properties 1, 2. These compounds exhibit increased structural anisotropy, dimensional changes, and the formation of novel structural types 3. The structural modifications induced by variations in bonding promote the emergence of new physical phenomena and offer opportunities to tune both optoelectronic and thermoelectronic properties by altering the electronic structure 4, 5, 6, 7, 8.
@@
CuBiSeCl2, a mixed-metal chalcohalide, stands out for its potential to exhibit both low thermal conductivity and favorable electronic properties. Previous experimental studies have shown that CuBiSeCl2 has a band gap of 1.33 eV, as measured by diffuse reflectance UV-vis spectroscopy, making it suitable for photovoltaic applications 3. Additionally, it demonstrates a low thermal conductivity of 0.27 W m-1 K-1 at room temperature 3. However, the complex crystal structure has limited comprehensive studies on the underlying mechanisms responsible for its low $\kappa_{L}$ , particularly for such intricate mixed-anion compounds.
@@
In this study, we investigate the microscopic mechanisms underlying the low $\kappa_{L}$ of CuBiSeCl2 by combining machine learning potentials with anharmonic lattice dynamics. We integrate machine learning potentials with the Boltzmann transport equation (BTE) and molecular dynamics (MD) simulations to conduct a detailed analysis of phonon transport and scattering mechanisms. This computational framework allows for more accurate predictions of thermal conductivity and provides deeper insights into the phonon anharmonicity and glassy thermal transport in crystalline compounds. Our results not only elucidate the origins of the low $\kappa_{L}$ in CuBiSeCl2 but also offer a new strategy for further reducing $\kappa_{L}$ through the delocalization of atoms and the introduction of softer chemical and vibrational bonds. The computational details can be found in the Supplemental Material.
Results and discussion
A. Crystal structure
@@
The crystal structure of CuBiSeCl2 belongs to space group $Pnma$ with number of 62 and the optimized lattice constants, using PBE potentials, are $a$ = 8.73 Å, $b$ = 4.05 Å, and $c$ = 12.99 Å, in which the van der Waals (vdW) interactions are described by the optB86b functional 9, 10. These values are consistent with the experimental values, which are $a$ = 8.78 Å, $b$ = 4.00 Å, and $c$ = 13.14 Å 3. As illustrated in Fig. 1(b), the phonon spectrum of the PBE functional is free of imaginary frequencies, indicating the dynamical stability of CuBiSeCl2. The primitive cell comprises 20 atoms: four Bi, four Se, four Cu, and eight Cl respectively.
B. Anharmonic lattice dynamics
The impact of fourth-order anharmonicity on the phonon spectra at finite temperatures is depicted in Fig. 1(b), using the interatomic force constants (IFCs) obtained from MLP. The overall phonon hardening is observed in crystalline CuBiSeCl2 as the temperature increases from 0 to 700 K. This suggests the presence of strong lattice anharmonicity in CuBiSeCl2 and highlights the importance of properly treating temperature renormalization 11, 12, 13, 14, 15. In particular, the low-frequency phonons below 3 Thz are predominantly influenced by the Cu (yellow) and Bi (red) atoms, as evidenced by the projected phonon density of states (PDOS) in Fig. 1(c).
Furthermore, atom-decomposed partial DOS in Fig. 1(c) and atomic participation ratios in Figs. 1(d-e) show that Cu and Bi atoms contribute mostly to the low-lying flattened phonon bands in the frequency lower than 2.0 THz. This suggests that strong phonon scattering is occurring, which suppresses the primary heat carriers, i.e., acoustic phonon modes 16, 17, 18. In contrast, the high-frequency phonon modes around 6 THz in CuBiSeCl2 are predominantly contributed by Se and Cl atoms and exhibit a temperature-dependent softening in Fig. 1(c). Cumulated $\kappa_{p}$ shown in Fig. 1(f) also approve that low frequency phonons less than 1.5 THz are the main contributor to the $\kappa_{p}$ of CuBiSeCl2.
C. Delocalization of copper atoms
Fig. 2(a) shows the mean square displacements (MSDs) of atoms in CuBiSeCl2 as a function of temperature along the $b$-axis. MSD is a measure of the deviation of the position of an atom with respect to a reference position over time. It indicates the atomic displacements of Cu are significantly larger than those of other atoms Bi, Se, and Cl, suggesting that the bonding of Cu atoms is heavily delocalized and soft 11.
In Fig. 2(b), the displacement of Cu atoms shows a relatively deep energy well with a flat bottom as a function of the phonon mode along the $b$-axis. The U-shaped potential energy curve significantly deviates from the harmonic approximation (quadratic fit) shown as the dotted line in Fig. 2(b). The U-shaped potential energy of Cu atom displacement can be well fitted by a fourth-order polynomial (solid yellow line in Fig. 2(b)), expressed as $y=1.41x^{4}+0.38x^{2}$, indicating strong quartic phonon anharmonicity.
The red color corresponds to complete localization and blue corresponds to complete delocalization of electrons in 2D electron localization function (ELF) of CuBiSeCl2 shown in Fig. 2(c). It can be seen that the localized electronic density around Cu is extremely low, significantly lower than the other atoms, which reveals that the Se and Cl elements exhibit a much stronger attraction of electrons than the Cu elements, suggesting a greater degree of delocalization for Cu atoms.
The $\kappa_{L}$ of CuBiSeCl2 at 300 K in Fig. 2(d-f), calculated using homogeneous nonequilibrium molecular dynamics (HNEMD) is the averaged result from five 2 ns simulation runs, totally 10 ns 19, 20, 21. We employed precise NEP to obtain the intrinsic $\kappa_{L}$ of CuBiSeCl2, highlighted in red color in Fig.2(d-f). The thicker solid line corresponds to the averaged result, and the converged $\kappa_{L}$ values are marked on the right side of the figure. These values indicate a significantly lower intrinsic $\kappa_{L}$ along the $c$-axis (0.34 W m-1 K-1) compared to the $a$- and $b$-axes (0.50 and 0.55 W m-1 K-1), consistent with the final result shown in Fig. 4(d) 13, 22.
Based on Fig. 2(a-c), we observe a significant difference in the delocalization of Cu compared to Bi, Se, and Cl. To investigate the impact of the delocalization of different atoms on $\kappa_{L}$, we used the same HNEMD method to calculate $\kappa_{L}$ along $a$-, $b$- and $c$-axes of the system after fixing different types of atoms. The results are shown in Fig. 2(d-f). Yellow, purple, and blue represent the fixed atoms of Cu, Bi, and Se, respectively. In all three directions, for the cases where Bi and Se atoms are fixed, a significant decrease in $\kappa_{L}$ compared to the intrinsic values is observed, indicating that the strong localization of Bi and Se atoms enhances the phonon transport of CuBiSeCl2. However, fixing Cu in the $a$- and $b$-axes reveals its counterintuitive phenomenon, as its fixation leads to an increase in $\kappa_{L}$ compared to the intrinsic value by 0.08 and 0.67 W m-1 K-1, respectively. This highlights that the delocalization of Cu has a negative impact on $\kappa_{L}$, especially on the $b$-axis 23.
Figure 2: (a) The mean square atomic displacements (MSDs) with temperatures for CuBiSeCl2. (b) Potential energy per atom with displacement of Cu (yellow) atoms along the $b$-axis direction with quartic and quadratic polynomial fitting. It writes as $y=1.41x^{4}+0.38x^{2}$ and $y=1.58x^{2}$. For the Bi (red), Se (blue), and Cl (purple) atoms, the quartic polynomial fittings along the $b$-axis direction are as follows: $y=0.67x^{4}+1.75x^{2}$, $y=1.20x^{4}+1.39x^{2}$ and $y=0.015x^{4}+2.76x^{2}$. The cubic term is minimal and less than $10^{-5}$ 24, 22. (c) Electron localization function (ELF) of CuBiSeCl2. The red color corresponds to complete localization and the blue color corresponds to complete delocalization of electrons. $\kappa_{L}$ predicted by the HNEMD method with NEP vs integration time for CuBiSeCl2 in $a$-axis (d), $b$-axis (e), and $c$-axis (f) respectively. The values labeled on the right represent the averaged final converged $\kappa_{L}$.
D. Thermal transport properties
@@
Based on Peierls-Boltzmann transport theory, the lattice thermal conductivity $\kappa_{p}$ is given by,
@@
and $\kappa_{c}$ describing tunneling and wavelike of coherent phonons, governed by the Wigner transport equation, is defined as 25, 13,
@@
The total $\kappa_{L}$ = $\kappa_{p}$ + $\kappa_{c}$, and the parameters in these equations can be found in the Supplementary Material.

### v3

@@ Introduction
Mixed-anion compounds, such as halide-chalcogenides, have recently gained attention as promising candidates for thermoelectric and photovoltaic applications due to their ability to achieve low thermal conductivity while maintaining favorable electronic properties 1, 2, 3. These compounds exhibit increased structural anisotropy, dimensional changes, and the formation of novel structural types 4. The structural modifications induced by variations in bonding promote the emergence of new physical phenomena and offer opportunities to tune both optoelectronic and thermoelectronic properties by altering the electronic structure 5, 6, 7, 8, 9.
@@
CuBiSeCl2, a mixed-metal chalcohalide, stands out for its potential to exhibit both low thermal conductivity and favorable electronic properties. Previous experimental studies have shown that CuBiSeCl2 has a band gap of 1.33 eV, as measured by diffuse reflectance UV-vis spectroscopy, making it suitable for photovoltaic applications 4. Additionally, it demonstrates a low thermal conductivity of 0.27 W m-1 K-1 at room temperature 4. However, the complex crystal structure has limited comprehensive studies on the underlying mechanisms responsible for its low $\kappa_{L}$ , particularly for such intricate mixed-anion compounds.
@@
In this study, we investigate the microscopic mechanisms underlying the low $\kappa_{L}$ of CuBiSeCl2 by combining machine learning potentials with anharmonic lattice dynamics. We integrate machine learning potentials with the Boltzmann transport equation (BTE) and molecular dynamics (MD) simulations to conduct a detailed analysis of phonon transport and scattering mechanisms. This computational framework allows for more accurate predictions of thermal conductivity and provides deeper insights into the phonon anharmonicity and glassy thermal transport in crystalline compounds. Our results not only elucidate the origins of the low $\kappa_{L}$ in CuBiSeCl2 but also offer a new strategy for further reducing $\kappa_{L}$ through the delocalization of atoms and the introduction of softer chemical and vibrational bonds.
Computational Methods
A. Structural optimization
The Vienna Ab initio Simulation Package (VASP) was used to optimize the crystal structure of CuBiSeCl2 10. All first-principles calculations are performed by using the PAW method with PBE exchange-correlation functional 11. The cut-off energy of 400 eV was used for the wave functions. The van der Waals (vdW) interactions are described by using optB86b functional 12, 13. The force and energy convergence thresholds of 10-5 eV/Å  and 10-10 eV respectively, were used for both structural relaxation and self-consistent density functional theory (DFT) calculations. The cell and the atoms were fully relaxed in the structural optimization process with the Gamma-centered k-point grids of 5×10×3.
@@ A. Structural optimization
The crystal structure of CuBiSeCl2 belongs to space group $Pnma$ with number of 62 and the optimized lattice constants, using PBE potentials, are $a$ = 8.73 Å, $b$ = 4.05 Å, and $c$ = 12.99 Å, in which the van der Waals (vdW) interactions are described by the optB86b functional 12, 13. These values are consistent with the experimental values, which are $a$ = 8.78 Å, $b$ = 4.00 Å, and $c$ = 13.14 Å 4. As illustrated in Fig. 1(b), the phonon spectrum of the PBE functional is free of imaginary frequencies, indicating the dynamical stability of CuBiSeCl2. The primitive cell comprises 20 atoms: four Bi, four Se, four Cu, and eight Cl respectively.
B. Data Generation
The dataset was generated using random perturbations with the dpdata package 14, 15. The deformation ratio of the unit cell was set to 4%, with perturbations in the diagonal part uniformly distributed within the range [-4%, 4%], and perturbations in the off-diagonal part uniformly distributed within the range [-2%, 2%]. The maximum atomic displacement radius was set to 0.4 Å, meaning that the atomic positions were randomly and uniformly distributed within a spherical space of radius 0.4 Å. A total of 600 structures were generated, of which 1/12 were randomly selected as the validation set, while the remaining were used as the training set. The structures of CuBiSeCl2 in the dataset were expanded by a 2×4×2 supercell which contains 320 atoms, and the k-point sampling was set to 1×1×1. The van der Waals (vdW) interactions is also included.
C. Training machine learning potential
To describe the interatomic interaction, machine learning neuroevolution potentials (NEP) was trained on the dataset generated by the DFT calculation 16, 17.
The NEP parameters can be fitted by minimizing the function 16,
$L(z)=$ $\lambda_{e}\left(\frac{1}{N_{\text{str}}}\sum_{n=1}^{N_{\text{str}}}\left(U^{\text{NEP}}(n,z)-U^{\text{tar}}(n)\right)^{2}\right)^{1/2}$ (1)
$+\lambda_{f}\left(\frac{1}{3N}\sum_{i=1}^{N}\left(F_{i}^{\text{NEP}}(z)-F_{i}^{\text{tar}}\right)^{2}\right)^{1/2}$
$+\lambda_{v}\left(\frac{1}{6N_{\text{str}}}\sum_{n=1}^{N_{\text{str}}}\sum_{\mu\nu}\left(W_{\mu\nu}^{\text{NEP}}(n,z)-W_{\mu\nu}^{\text{tar}}(n)\right)^{2}\right)^{1/2}$
$+\lambda_{1}\frac{1}{N_{\text{par}}}\sum_{n=1}^{N_{\text{par}}}|z_{n}|$
$+\lambda_{2}\left(\frac{1}{N_{\text{par}}}\sum_{n=1}^{N_{\text{par}}}z_{n}^{2}\right)^{1/2},$
where $N_{\text{str}}$ is the number of structures in the training dataset and $N$ is the total number of atoms in all these structures. The first three terms are the weighted averages of the energy, forces, and virial tensor using NEP compared with the DFT results. That is, the loss functions for energy, force, and virial are defined as their root mean square errors (RMSEs) between the current NEP predictions and the target values. The last two terms represent $\ell_{1}$ and $\ell_{2}$ regularization. The weights $\lambda_{e}$, $\lambda_{f}$, $\lambda_{v}$, $\lambda_{1}$, and $\lambda_{2}$ are tunable hyperparameters. When calculating the loss function, we use the following units: eV/atom for energy and virial, and eV/Å for force components.
Since the cutoff values for cubic and quartic interactions used in the phonon thermal conductivity calculations are 4.12 Å and 3.30 Å, both smaller than the NEP fitting radial cutoff radius 7.0 Å and angular cutoff radius 7.0 Å, indicating the NEP fitting cutoff radius is sufficiently large for the thermal conductivity calculation.
Furthermore, we set the number of basis functions that are used to build the radial and angular descriptor to 14, which is greater than the default value of 8, ensuring sufficient accuracy.
D. Phonon scattering properties
Three-phonon scattering rates can be expressed as:
$\Gamma^{+}_{\lambda\lambda^{\prime}\lambda^{\prime\prime}}=$ $\frac{\hbar\pi}{4}\frac{f_{0}^{\prime}-f_{0}^{\prime\prime}}{\omega_{\lambda}\omega_{\lambda^{\prime}}\omega_{\lambda^{\prime\prime}}}\lvert V^{+}_{\lambda\lambda^{\prime}\lambda^{\prime\prime}}\rvert^{2}$ (2)
$\delta\left(\omega_{\lambda}+\omega_{\lambda^{\prime}}-\omega_{\lambda^{\prime\prime}}\right)$
$\Gamma^{-}_{\lambda\lambda^{\prime}\lambda^{\prime\prime}}=$ $\frac{\hbar\pi}{4}\frac{f_{0}^{\prime}+f_{0}^{\prime\prime}+1}{\omega_{\lambda}\omega_{\lambda^{\prime}}\omega_{\lambda^{\prime\prime}}}\lvert V^{-}_{\lambda\lambda^{\prime}\lambda^{\prime\prime}}\rvert^{2}$ (3)
$\delta\left(\omega_{\lambda}-\omega_{\lambda^{\prime}}-\omega_{\lambda^{\prime\prime}}\right)$
where, for simplicity, $f_{0}^{\prime}$ stands for $f_{0}\bigl{(}\omega_{\lambda^{\prime}}\bigr{)}$ and $f_{0}^{\prime\prime}$ denotes $f_{0}\bigl{(}\omega_{\lambda^{\prime\prime}}\bigr{)}$. The term $\Gamma_{\lambda\lambda^{\prime},\lambda^{\prime\prime}}^{+}$ corresponds to an absorption process, where two incident phonons $(\omega_{\lambda},\omega_{\lambda^{\prime}})$ merge into a single phonon whose frequency is their sum. Meanwhile, $\Gamma_{\lambda\lambda^{\prime},\lambda^{\prime\prime}}^{-}$ describes an emission process, in which one phonon splits into two $(\omega_{\lambda^{\prime}},\omega_{\lambda^{\prime\prime}})$ 18.
Phonon density of states (DOS) is calculated with a linear tetrahedron method which is default in Phonopy 19. Projected DOS is calculated using “PDOS = AUTO”. Without specifying projection direction, PDOS is computed as sum of $g^{j}(\omega,\hat{\mathbf{n}})$ projected onto cartesian axes $x,y,z$,
$g^{j}(\omega)=\sum_{\hat{\mathbf{n}}=\{x,y,z\}}g^{j}(\omega,\hat{\mathbf{n}}).$ (4)
where $j$ is the atom indices and $\hat{\mathbf{n}}$ is the unit projection direction vector.
Mean square displacement of different atoms is also calculated using Phonopy package 19. The expectation value of the squared atomic displacement is calculated as,
$\left\langle|u^{\alpha}(jl,t)|^{2}\right\rangle=$ $\frac{\hbar}{2Nm_{j}}\sum_{\mathbf{q},\nu}\omega_{\nu}(\mathbf{q})^{-1}$ (5)
$\left(1+2n_{\nu}(\mathbf{q},T)\right)|e_{\nu}^{\alpha}(j,\mathbf{q})|^{2},$
where $n_{\nu}(\mathbf{q},T)$ is the phonon population, which is given by,
$n_{\nu}(\mathbf{q},T)=\frac{1}{\exp(\hbar\omega_{\nu}(\mathbf{q})/k_{\text{B}}T)-1},$ (6)
$T$ is the temperature, and $k_{\text{B}}$ is the Boltzmann constant.
Atomic participation ratio (APR) is calculated with alamode package and is defined as follow:
$APR_{\mathbf{q}j,\kappa}$ $=\frac{|\mathbf{e}(\kappa;\mathbf{q}j)|^{2}}{M_{\kappa}}$ (7)
$\quad/\left(N_{\kappa}\sum_{\kappa}^{N_{\kappa}}\frac{|\mathbf{e}(\kappa;\mathbf{q}j)|^{4}}{M_{\kappa}^{2}}\right)^{1/2}$
APR is an atomic decomposition of PR that satisfies $PR_{\mathbf{q}j}=\sum_{\kappa}(APR_{\mathbf{q}j,\kappa})^{2}$.
E. Lattice thermal conductivity
@@
The lattice thermal conductivity $\kappa_{p}$ based on Peierls-Boltzmann transport theory is given by20,
@@
where $\hbar$ denotes the reduced Planck constant, $k_{B}$ the Boltzmann constant, $T$ the absolute temperature, $V$ the volume of the primitive unit cell, and $N_{0}$ the number of phonon wave vectors sampled in the first Brillouin zone. The variables $n_{\lambda}$, $\omega_{\lambda}$, $\bm{v}_{\lambda}$, and $\tau_{\lambda}$ indicate the phonon population, frequency, group velocity, and lifetime for the $\lambda$ mode which is defined by wave vector $q$ and branch index $s$.
In the interatomic force constants (IFCs) calculations, a 2$\times$4$\times$1 supercell of CuBiSeCl2 was used. Third- and fourth-order IFCs were calculated up to the eighth and third nearest neighbors, respectively 21, 22. The q-mesh grid was set to 6$\times$12$\times$4 for four-phonon (4ph) processes and 10$\times$20$\times$7 for three-phonon (3ph) processes. The self-consistent phonon (SCPH) approximation 23, 24 was applied to incorporate first-order phonon frequency corrections due to quartic anharmonicity, which enhances accuracy for materials with soft phonon modes and significant anharmonicity. The SCPH is described by
${\Omega}_{\lambda}^{2}={\omega}_{\lambda}^{2}+2{\Omega}_{\lambda}\sum\limits_{{\lambda}_{1}}I_{\lambda\lambda_{1}},$ (9)
where $\omega_{\lambda}$ represents the initial phonon frequency derived from the harmonic approximation, and $\Omega_{\lambda}$ denotes the temperature-dependent renormalized phonon frequency. The scalar $I_{\lambda\lambda_{1}}$ is calculated as follows:
${I_{\lambda\lambda_{1}}}=\frac{\hbar}{8N_{0}}\frac{V^{(4)}(\lambda,-\lambda,\lambda_{1},-\lambda_{1})}{\Omega_{\lambda}\Omega_{\lambda_{1}}}\left[1+2n_{\lambda}(\Omega_{\lambda_{1}})\right],$ (10)
where $V^{(4)}$ represents the fourth-order IFCs in the reciprocal space, and the phonon population $n_{\lambda}$ follows the Bose-Einstein distribution. Eq. (9) and Eq. (10) both possess the parameters $I_{\lambda\lambda_{1}}$ and $\Omega_{\lambda}$, allowing the SCPH equation to be solved iteratively. Also, we fix the volume in self-consistent phonon calculation which means that thermal expansion was not considered in temperature-dependent renormalized phonon frequency 25.
To enhance the accuracy of the lattice thermal conductivity, the coherent phonon $\kappa_{c}$ describing tunneling of coherent phonons, was also included and is defined as26
@@
where $\hbar$ is the reduced Planck constant, $k_{B}$ is Boltzmann constant, $V$ is the primitive cell volume, and $N_{0}$ is the total number of sampled phonon vectors.
Additionally, the temperature-dependent $\kappa_{L}$ of CuBiSeCl2 is also calculated from homogeneous nonequilibrium molecular dynamics (HNEMD) simulations 27, 28, 29, an average result of 2ns for 5 times with the model cell around 75 Å along each edge. The total heat flow in the system is the sum of contributions from each atom,
$J=\sum_{i}\sum_{j\neq i}\mathbf{r}_{ij}\left(\frac{\partial U_{j}}{\partial\mathbf{r}_{ji}}\cdot\mathbf{v}_{i}\right)$ (12)
where J in the Green–Kubo method for calculating thermal conductivity, with ${r}_{ij}$ is the distance vector between atoms i and j, U is the potential energy, and v is velocity. In the HNEMD method, each atom in the system is subjected to an external force to drive it out of equilibrium. This force, expressed as
$\vec{F}_{\text{ext},i}=E_{i}\vec{F}_{e}+\vec{F}_{e}\cdot\vec{W}_{i},$ (13)
This approach offers a way to compute thermal conductivity. Unlike the NEMD method, the HNEMD framework maintains the system in a uniform non-equilibrium state, without explicit heat sources or sinks. As a result, phonon boundary scattering is absent, and heat circulates due to periodic boundary conditions along the transport direction. This method produces finite-size effects similar to those seen in the EMD approach.
Results and discussion
A. Anharmonic lattice dynamics
The impact of fourth-order anharmonicity on the phonon spectra at finite temperatures is depicted in Fig. 1(b), using the interatomic force constants (IFCs) obtained from MLP. The overall phonon hardening is observed in crystalline CuBiSeCl2 as the temperature increases from 0 to 700 K. This suggests the presence of strong lattice anharmonicity in CuBiSeCl2 and highlights the importance of properly treating temperature renormalization 30, 23, 31, 32, 33. In particular, the low-frequency phonons below 3 Thz are predominantly influenced by the Cu (yellow) and Bi (red) atoms, as evidenced by the projected phonon density of states (PDOS) in Fig. 1(c).
Furthermore, atom-decomposed partial DOS in Fig. 1(c) and atomic participation ratios in Figs. 1(d-e) show that Cu and Bi atoms contribute mostly to the low-lying flattened phonon bands in the frequency lower than 2.0 THz.
In contrast, the high-frequency phonon modes around 6 THz in CuBiSeCl2 are predominantly contributed by Se and Cl atoms and exhibit a temperature-dependent softening in Fig. 1(c). Cumulated $\kappa_{p}$ shown in Fig. 1(f) also approve that low frequency phonons less than 1.5 THz are the main contributor to the $\kappa_{p}$ of CuBiSeCl2. These indicate that the phonon vibrations of Cu and Bi atoms predominantly govern $\kappa_{p}$ in CuBiSeCl234, 35, 36.
B. Delocalization of copper atoms
Fig. 2(a) shows the mean square displacements (MSDs) of atoms obtained by both harmonic approximation (HA) and self-consistent phonon (SCPH) in CuBiSeCl2 as a function of temperature calculations along the $b$-axis. MSD is a measure of the deviation of the position of an atom with respect to a reference position over time. It indicates the atomic displacements of Cu are significantly larger than those of other atoms Bi, Se, and Cl, suggesting that the bonding of Cu atoms is heavily delocalized and soft 30.
Meanwhile, the MSDs calculated by HA typically exhibits overestimation, while anharmonic effects suppress the vibrational amplitudes of all atoms 30. The discrepancy between them can intuitively quantify the role of anharmonic effects. Fig. 2(b) illustrates the differences in MSDs between the HA and SCPH calculations along all three axes for CuBiSeCl2. It can be observed that the MSDs differences of Cu atoms are larger than those of other elements in all three directions. Particularly, the MSDs difference along $b$-axis is the largest, indicating strong phonon anharmonicity in this direction. Besides, the U-shaped potential energy of Cu atom displacement can be well fitted by a fourth-order polynomial in Fig. S8 37, also qualitatively indicating strong high-order phonon anharmonicity 38.
The red color corresponds to complete localization and blue corresponds to complete delocalization of electrons in 2D electron localization function (ELF) of CuBiSeCl2 shown in Fig. 2(c). The deeper blue color between Cu and Cl or Se atoms indicates there is no significant electron wave function overlap between Cu and its surrounding atoms, further suggesting that no strong bonding is formed and a greater degree of delocalization for Cu atoms.
The $\kappa_{L}$ of CuBiSeCl2 at 300 K in Fig. 2(d-f), calculated using homogeneous nonequilibrium molecular dynamics (HNEMD) is the averaged result from five 2 ns simulation runs, totally 10 ns 27, 28, 29. We employed precise NEP to obtain the intrinsic $\kappa_{L}$ of CuBiSeCl2, highlighted in red color in Fig.2(d-f). The thicker solid line corresponds to the averaged result, and the converged $\kappa_{L}$ values are marked on the right side of the figure. These values indicate a significantly lower intrinsic $\kappa_{L}$ along the $c$-axis (0.34 W m-1 K-1) compared to the $a$- and $b$-axes (0.50 and 0.55 W m-1 K-1), consistent with the final result shown in Fig. 4(d) 31, 39.
Based on Fig. 2(a-c), we observe a significant difference in the delocalization of Cu compared to Bi, Se, and Cl. To investigate the impact of the delocalization of different atoms on $\kappa_{L}$, we used the same HNEMD method to calculate $\kappa_{L}$ along $a$-, $b$- and $c$-axes of the system after fixing different types of atoms by setting the velocities and forces acting on the selected atoms to zero with the function implanted in GPUMD software package 16. The results are shown in Fig. 2(d-f). Yellow, purple, and blue represent the fixed atoms of Cu, Bi, and Se, respectively. In all three directions, for the cases where Bi and Se atoms are fixed, a significant decrease in $\kappa_{L}$ compared to the intrinsic values is observed, indicating that the strong localization of Bi and Se atoms enhances the phonon transport of CuBiSeCl2. However, fixing Cu in the $a$- and $b$-axes reveals its counterintuitive phenomenon, as its fixation leads to an increase in $\kappa_{L}$ compared to the intrinsic value by 0.08 and 0.67 W m-1 K-1, respectively. This highlights that the delocalization of Cu has a negative impact on $\kappa_{L}$, especially on the $b$-axis 40.
Figure 2: (a) The mean square atomic displacements (MSDs) with temperatures for CuBiSeCl2. The solid and dashed lines correspond to the MSDs obtained by harmonic approximation (HA) and self-consistent phonon (SCPH) calculations respectively. (b) Difference between MSDs obtained by HA and SCPH approaches in $a$-axis, $b$-axis and $c$-axis. (c) Electron localization function (ELF) of CuBiSeCl2. The red color corresponds to complete localization and the blue color corresponds to complete delocalization of electrons. $\kappa_{L}$ predicted by the HNEMD method with NEP vs integration time for CuBiSeCl2 in $a$-axis (d), $b$-axis (e), and $c$-axis (f) respectively. The values labeled on the right represent the averaged final converged $\kappa_{L}$.
C. Thermal transport properties